Please help with this problem:Becoming a millionaire: How much would you have to invest now at 6.5% interest compounded semi-annually to have a million dollars in 40 years?: Please help with this problem:Becoming a millionaire: How much would you have to invest now at 6.5% interest compounded semi-annually to have a million dollars in 40 years?
Answer by checkley77(1718)

1,000,000=x(1+.065/2)^40*2
1,000,000=x(1.0325)^80
1,000,000=12.9183x
x=1,000,000/12.9183
x=$77,409.57 would be the initial investment.